• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name Z. Chai

    Log entry time 08:06:20 on June17,2000

    Entry number 44011

    keyword=Owl shift summary



    00:00 Start owl shift. Continue taking data for kin(135,180).
    Beam current 93uA.
    02:58 Beam tripped because of the RF problem in north linac.
    03:30 MCC told us they have some HV problem and the on-call person
    is on the way. It may take one hour to fix that problem.
    04:30 The target coolan pressure is fluctuating and increasing to the
    set HIGH limit.
    See halog 43999.
    05:15 Beam back to Hall A. Ask for 50 uA beam current to avoid the overpressure of the coolant.
    Tried to increase the coolant back to its original flow rate, its pressure
    also went up. Since the target oncall doesn't know how to get rid
    of this coolant pressure problem and MCC can't solve this probelm for us either,
    we just keep data taking at a lower beam current (50uA) so that the coolant flow can
    be set to 30 and the coolant pressure is not increasing any more. The experts should
    be called to check this problem during the day.
    06:57 Open the JT valve to 33, its pressure looks stable now (54.6). Ask MCC
    to increase current to 60uA.
    07:45 Ask MCC to increase current to 70uA since the coolant pressure is stable
    at this flow rate.

    Time Run Events Comment
    23:47 2153 1.3M 93uA
    00:04 2154 1.3M
    00:21 2155 1.3M
    00:39 2156 1.3M
    00:57 2157 1.3M
    01:12 2158 1.3M
    01:32 2159 1.3M
    01:50 2160 1.3M
    02:08 2161 1.3M
    02:25 2162 1.3M 90uA
    02:43 2163 0.5M RF problem in north linac
    05:15 2164 1.3M I_beam=50uA
    05:57 2165 1.3M 50uA
    06:39 2166 1.3M Beam current is increased to 60uA at 6:57.
    07:08 2167 1.3M 60uA, at 7:45 increased to 70uA
    07:47 2068